# Beatty Sequences for a Quadratic Irrational: Decidability and Applications

### 24 Mar 2024 v2

@@ Abstract
Let $\alpha$ and $\beta$ belong to the same quadratic field. We show that the inhomogeneous Beatty sequence $(\lfloor n\alpha+\beta\rfloor)_{n\geq 1}$ is synchronized, in the sense that there is a finite automaton that takes as input the Ostrowski representations of $n$ and $y$ in parallel, and accepts if and only if $y=\lfloor n\alpha+\beta\rfloor$. Since it is already known that the addition relation is computable for Ostrowski representations based on a quadratic number, a consequence is a new and rather simple proof that the first-order logical theory of these sequences with addition is decidable. The decision procedure is easily implemented in the free software Walnut.
@@ 1 Introduction
Let $\alpha$ be an irrational number. The sequence $(\lfloor n\alpha\rfloor)_{n\geq 1}$ is known as a Beatty sequence [6] and has been very widely studied. [^1] Some authors [14] insist that $\alpha$ must be greater than $1$, but we do not. Also, sometimes it is useful to allow the index $n$ to be 00. The two most famous Beatty sequences correspond to $\alpha=(1+\sqrt{5})/2$ and $\alpha=(3+\sqrt{5})/2$; they are also known as the lower Wythoff and upper Wythoff sequences, respectively, and they form sequences A000201 and A001950 in the On-Line Encyclopedia of Integer Sequences (OEIS) [46].
In an inhomogeneous Beatty sequence [^2] Sometimes called non-homogeneous Beatty sequence in the literature; see [17]., we permit a constant additive factor inside the floor brackets: $(\lfloor n\alpha+\beta\rfloor)_{n\geq 1}$. The first difference of such a sequence forms a Sturmian sequence, which also have been widely studied and have a huge literature [35, 9].
@@
In this paper, we show that if $\alpha$ and $\beta$ belong to the same quadratic field, then the inhomogeneous Beatty sequence $(\lfloor n\alpha+\beta\rfloor)_{n\geq 1}$ is synchronized, in the sense that there is a finite automaton that takes the Ostrowski representations of $n$ and $y$ in parallel, and accepts if and only if $y=\lfloor n\alpha+\beta\rfloor$. Ostrowski representations are discussed in Section 3; for an introduction and explanation of how automata can use them, see [4]. Synchronized sequences are discussed in Section 4; for a summary, see [43]. As a consequence, we get a new and rather simple proof that first-order logical theory of quadratic Beatty sequences with addition is decidable. History of this result is discussed in Section 2, and our results appear in Section 5. Not only are these first-order statements decidable, but in practice many theorems about individual Beatty sequences can be proved in a matter of seconds with an implementation of the decision procedure in the free software Walnut [36, 45]. Walnut is available at
@@
In Section 8 we use these ideas to solve some open problems of Don Reble. In Section 9 we give an explicit description of a sequence studied by Hildebrand, Li, Li, and Xie. In Sections 10 and 11, we solve some open problems of Kimberling. Finally, in Section 12, we prove a number of identities about sequences in the OEIS.
@@ 2 Previous work
As a consequence, the first-order theory of $\mathbb{N}$ together with the Sturmian sequence ${\bf f}=01001010\cdots$ (the Fibonacci word, corresponding to $(3-\sqrt{5})/2$) with addition is decidable. This was observed by Bruyère and Hansel [11], building on work of Bruyère et al. [12]. Later, these ideas were used by Mousavi et al. [37] to solve questions about the properties of $\bf f$ and related words. Also see [24].
@@ 3 Ostrowski $\alpha$-representation
In the famous Fibonacci (or Zeckendorf) system for representing natural numbers [34, 48], every integer is written uniquely as a sum of distinct Fibonacci numbers, subject to the condition that we never use two consecutive Fibonacci numbers. More precisely, if the Fibonacci numbers are (as usual) $F_{0}=0$, $F_{1}=1$, and $F_{n}=F_{n-1}+F_{n-2}$ for $n\geq 2$, then a natural number $n$ can be expressed as a binary string $e_{t}\cdots e_{1}$ with interpretation $n=\sum_{1\leq j\leq t}e_{i}F_{i+1}$. Notice that the most significant digit is written at the left.
The Ostrowski numeration system is a generalization of this idea. Let $\alpha>0$ be an irrational real number, with continued fraction $[a_{0},a_{1},a_{2},\ldots]$. Define, as usual, the $i$’th convergent $p_{i}/q_{i}$ to be $[a_{0},a_{1},\ldots,a_{i}]$. Recall (e.g., [23]) that $p_{i}=a_{i}p_{i-1}+p_{i-2}$ and $q_{i}=a_{i}q_{i-1}+q_{i-2}$.
@@
See [40] and [2, §3.9]. We write the canonical Ostrowski $\alpha$-representation of $n$, starting with the most significant digit, and satisfying the rules (a)–(c) above, as $(n)_{\alpha}$. The inverse mapping, which sends a string $x=e_{t}e_{t-1}\cdots e_{0}$ to the sum $\sum_{0\leq i\leq t}e_{i}q_{i}$, is denoted by $[x]_{\alpha}$.
In some cases we need to write the $\alpha$-representation of pairs or tuples. We do so by working over a product alphabet, and padding the shorter representation at the front with 00’s, if needed.
@@ 4 Automata, synchronized sequences, and first-order logic
We say that a sequence $(s_{n})_{n\geq 0}$ is synchronized if there is a finite automaton that takes the representations of $n$ and $y$ in parallel, and accepts if and only if $y=s_{n}$. In this paper, the particular representations we use are the Ostrowski $\alpha$-representations discussed in the previous section.
Synchronized sequences have a number of nice closure properties. These can be found in [43] for automatic sequences, but the same proofs also suffice for sequences defined in terms of Ostrowski representations (provided one knows that the addition relation is synchronized, as proved in [4]).
@@
    The sequence $(\lfloor(bn+c)/d\rfloor)$ is synchronized for integers $b,c,d$ with $b\geq\max(1,-c)$ and $d\geq 1$.
@@
    If $(f(n))_{n\geq 1}$ and $(g(n))_{n\geq 1}$ are both synchronized then so is their composition $f(g(n))_{n\geq 1}$.
@@
There is an algorithm that takes as input a first-order logical formula on natural numbers for $f(n)$, defined in terms of addition, subtraction, multiplication by constants, and synchronized sequences and computes a synchronized automaton for $f$.
@@ 5 The main result and its implementation
Suppose $10<\gamma<$ 1 is a quadratic irrational with purely periodic continued fraction
$[0,a_{1},a_{2},\ldots,a_{m},a_{1},a_{2},\ldots,a_{m},a_{1},a_{2},\ldots,a_{m},\ldots].$
Let $q_{i}$ denote the denominator of the $i$’th convergent to this continued fraction. Then for all integers $n\geq 1$ we have
[…]
@@
Since $a_{i}=a_{i+m}$ for all $i\geq 1$, we have
$\left[\begin{array}[]{cc}p_{i+m}&p_{i+m-1}\\ q_{i+m}&q_{i+m-1}\end{array}\right]$ $=\left[\begin{array}[]{cc}a_{0}&1\\ 1&0\end{array}\right]\left[\begin{array}[]{cc}a_{1}&1\\ 1&0\end{array}\right]\cdots\left[\begin{array}[]{cc}a_{i+m}&1\\ 1&0\end{array}\right]$
$=\left[\begin{array}[]{cc}a_{0}&1\\ 1&0\end{array}\right]\left[\begin{array}[]{cc}a_{1}&1\\ 1&0\end{array}\right]\cdots\left[\begin{array}[]{cc}a_{m}&1\\ 1&0\end{array}\right]\quad\left[\begin{array}[]{cc}a_{m+1}&1\\ 1&0\end{array}\right]\cdots\left[\begin{array}[]{cc}a_{i+m}&1\\ 1&0\end{array}\right]$
$=\left[\begin{array}[]{cc}a_{0}&1\\ 1&0\end{array}\right]\left[\begin{array}[]{cc}a_{1}&1\\ 1&0\end{array}\right]\cdots\left[\begin{array}[]{cc}a_{m}&1\\ 1&0\end{array}\right]\quad\left[\begin{array}[]{cc}a_{1}&1\\ 1&0\end{array}\right]\cdots\left[\begin{array}[]{cc}a_{i}&1\\ 1&0\end{array}\right]$
$=\left[\begin{array}[]{cc}p_{m}&p_{m-1}\\ q_{m}&q_{m-1}\end{array}\right]\left[\begin{array}[]{cc}q_{i}&q_{i-1}\\ p_{i}&p_{i-1}\end{array}\right],$
@@
$q_{i+m}=q_{m}q_{i}+q_{m-1}p_{i}.$ (3)
Now suppose the Ostrowski $\gamma$-representation of $n-1$ is $\sum_{0\leq i\leq t}e_{i}q_{i}$. Then from Eq. (3), we get
$[(n-1)_{\gamma}0^{m}]_{\gamma}$ $=\sum_{0\leq i\leq t}e_{i}q_{i+m}$
$=\sum_{0\leq i\leq t}e_{i}(q_{m}q_{i}+q_{m-1}p_{i})$
$=\left(q_{m}\sum_{0\leq i\leq t}e_{i}q_{i}\right)+\left(q_{m-1}\sum_{0\leq i\leq t}e_{i}p_{i}\right)$
$=q_{m}(n-1)+q_{m-1}\lfloor n\gamma\rfloor,$
@@
Suppose $10<\gamma<$ 1 is a quadratic irrational with purely periodic continued fraction. Then the sequence $(\lfloor n\gamma\rfloor)_{n\geq 1}$ is $\gamma$-Ostrowski synchronized.
@@
It is easy to see that for each $m\geq 1$ there is a finite automaton ${\tt shift}_{m}$ accepting $0^{m}y$ and $y0^{m}$ in parallel. We sketch the details. The states are $m$-tuples of integers appearing in the Ostrowski representation of $\mathbb{N}$, and the initial state and unique final state is $0^{m}$. Here the “meaning” of a state named $f_{1}\cdots f_{m}$ is that these are the last $m$ symbols of $y$ that the automaton has read. The transition function $\delta$ is defined by $\delta(f_{1}f_{2}\cdots f_{m},[f_{1},g])=f_{2}\cdots f_{m}g$.
Since $q_{m}$ and $q_{m-1}$ are constants, the first-order formula
$\exists u,v\ n=u+1\,\wedge\,{\tt shift}_{m}(u,v)\,\wedge\,v=q_{m-1}\cdot z+q_{m}\cdot u$
@@
Suppose $10<\gamma<$ 1 is a quadratic irrational with purely periodic continued fraction
$[0,a_{1},a_{2},\ldots,a_{m},a_{1},a_{2},\ldots,a_{m},a_{1},a_{2},\ldots,a_{m},\ldots],$
and suppose $\alpha,\beta\in\mathbb{Q}(\gamma)$ such that $\alpha\geq 0$ and $\alpha+\beta\geq 0$. Then $(\lfloor n\alpha+\beta\rfloor)_{n\geq 1}$ is synchronized in terms of Ostrowski $\gamma$-representations.
@@
Let $\gamma$ be a purely periodic quadratic irrational, and let $\alpha,\beta\in\mathbb{Q}(\gamma)$. The first-order theory of the natural numbers with addition and the function $n\rightarrow\lfloor n\alpha+\beta\rfloor$ is decidable.
@@
Next, we prove a sort of converse to Corollary 5. We need a lemma. Recall that the subword complexity (aka factor complexity) of a word is the function $\rho(n)$ counting the number of distinct length-$n$ factors.
Lemma 9.
Suppose $10<\alpha_{i},\beta_{i}<$ 1 are real numbers for $1\leq i\leq k$. Let
$1,\alpha_{1},\ldots,\alpha_{k}$
Proof.
By the multidimensional version of Kronecker’s theorem (see, e.g., [23, Thm. 442]), the linear independence criterion implies that there are integers $q$ and $p_{i}$, $1\leq i\leq k$ such that $|\sum_{1\leq i\leq k}q\alpha_{i}-p_{i}+b_{i}|<\epsilon$.
We can now prove a sort of converse to Corollary 5.
Theorem 10.
Suppose $10<\gamma<$ 1 is a quadratic irrational with purely periodic continued fraction expansion. Suppose the Beatty sequence $(\lfloor n\alpha+\beta\rfloor)_{n\geq 1}$ is $\gamma$-Ostrowski synchronized. Then $\alpha\in\mathbb{Q}(\gamma)$. Furthermore, the expression of $\alpha$ in terms of $\gamma$ is computable.
Proof.
Suppose $1,\alpha,\gamma$ are linearly independent. If the Beatty sequence $(\lfloor n\alpha+\beta\rfloor)_{n\geq 1}$ is $\gamma$-Ostrowski synchronized, then the corresponding Sturmian sequence (possibly recoded over the alphabet $\{0,1\}$ if necessary) is $\gamma$-Ostrowski automatic. This is also true for the Sturmian sequence associated with slope $\gamma$ and intercept 00. Then the product of these two sequences (over the alphabet $\{0,1\}^{2}$) is also $\gamma$-Ostrowski automatic. Hence this product has subword complexity $\Theta(n^{2})$, but we know that the subword complexity of an automatic sequence is $O(n)$. This contradiction shows that $1,\alpha,\gamma$ are linearly dependent, and hence there are integers $a,b,c$ such that $a\alpha=b\gamma+c$.
It remains to show that we can computably determine $a,b$, and $c$. To do so we iterate over all possible integer triples by increasing order of $|a|+|b|+|c|$. For each triple we construct the an $\gamma$-Ostrowski automaton for the Sturmian word with slope $(b\gamma+c)/a$. We then check (using the logical approach) whether the set of factors occurring in Sturmian word is the same as that for the Sturmian word with slope $\alpha$ and intercept $\beta$. Since all Sturmian words with the same slope have the same set of factors, this suffices. ∎

### 02 Apr 2026 v3

@@ Abstract
Let $\alpha$ and $\beta$ be real numbers belonging to the same quadratic field. We show that the inhomogeneous Beatty sequence $(\lfloor n\alpha+\beta\rfloor)_{n\geq 1}$ is synchronized, in the sense that there is a finite automaton that takes as input the Ostrowski representations of $n$ and $y$ in parallel, and accepts if and only if $y=\lfloor n\alpha+\beta\rfloor$. Since it is already known that the addition relation is computable for Ostrowski representations based on a quadratic irrational, a consequence is a new and rather simple proof that the first-order logical theory of these sequences with addition is decidable. The decision procedure is easily implemented in the free software Walnut.
@@ 1 Introduction
Let $\alpha$ be an irrational number. The sequence $(\lfloor n\alpha\rfloor)_{n\geq 1}$ is known as a Beatty sequence [6] and has been very widely studied. [^1] Some authors [14] insist that $\alpha$ must be greater than $1$, but here we do not. Also, sometimes it is useful to allow the index $n$ to be 0. The two most famous Beatty sequences correspond to $\alpha=(1+\sqrt{5})/2$ and $\alpha=(3+\sqrt{5})/2$; they are also known as the lower Wythoff and upper Wythoff sequences, respectively, and they form sequences A000201 and A001950 in the On-Line Encyclopedia of Integer Sequences (OEIS) [49].
In an inhomogeneous Beatty sequence [^2] Sometimes called non-homogeneous Beatty sequence in the literature; see [17]., we permit a constant additive factor inside the floor brackets: $(\lfloor n\alpha+\beta\rfloor)_{n\geq 1}$. The first difference of such a sequence forms a Sturmian sequence, which also has been widely studied with a huge literature [37, 9].
@@
In this paper, we show that if $\alpha$ and $\beta$ belong to the same quadratic field, then the inhomogeneous Beatty sequence $(\lfloor n\alpha+\beta\rfloor)_{n\geq 1}$ is synchronized, in the sense that there is a finite automaton that takes the Ostrowski representations of $n$ and $y$ in parallel, and accepts if and only if $y=\lfloor n\alpha+\beta\rfloor$. Furthermore, the converse of this theorem also holds. Ostrowski representations are discussed in Section 3; for an introduction and explanation of how automata can use them, see [4]. Synchronized sequences are discussed in Section 4; for a summary, see [46]. As a consequence, we get a new and rather simple proof that first-order logical theory of quadratic Beatty sequences with addition is decidable. History of this result is discussed in Section 2, and our principal results appear in Section 5: Theorem 9 and Corollary 11. Not only are these first-order statements decidable, but in practice many theorems about individual Beatty sequences can be proved in a matter of seconds with an implementation of the decision procedure in the free software Walnut [38, 48]. Walnut is available at
@@
In Section 8 we use these ideas to solve some open problems of Don Reble, and in Section 9 we show a question of Graham is decidable for quadratic irrationals. In Section 10 we give an explicit description of a sequence studied by Hildebrand, Li, Li, and Xie. In Sections 11 and 12, we solve some open problems of Kimberling. In Section 13, we prove a number of identities about sequences in the OEIS. In Section 14 we discuss fractional parts. Finally, in Section 15 we prove the converse (Theorem 28) of Theorem 9.
@@ 2 Previous work
As a consequence, the first-order theory of $\mathbb{N}=\{0,1,2,\ldots\}$ together with the Sturmian sequence ${\bf f}=01001010\cdots$ (the Fibonacci word, corresponding to $(3-\sqrt{5})/2$) with addition is decidable. This was observed by Bruyère and Hansel [11], building on work of Bruyère et al. [12]. Later, these ideas were used by Mousavi et al. [39] to solve questions about the properties of $\bf f$ and related words. Also see [25].
@@ 3 Ostrowski $\alpha$-representation
In the famous Fibonacci (or Zeckendorf) system for representing natural numbers [36, 51], every non-negative integer is written uniquely as a sum of distinct Fibonacci numbers, subject to the condition that we never use two consecutive Fibonacci numbers. More precisely, if the Fibonacci numbers are (as usual) $F_{0}=0$, $F_{1}=1$, and $F_{n}=F_{n-1}+F_{n-2}$ for $n\geq 2$, then a natural number $n$ can be expressed as a binary string $e_{t}\cdots e_{1}$ with interpretation $n=\sum_{1\leq j\leq t}e_{j}F_{j+1}$. Notice that the most significant digit is written at the left.
The Ostrowski numeration system is a generalization of this idea. Let $\alpha>0$ be an irrational real number, with continued fraction $a_{0},a_{1},a_{2},\ldots]$. Define, as usual, the $i$’th convergent $p_{i}/q_{i}$ to be $a_{0},a_{1},\ldots,a_{i}]$. Recall (e.g., [24]) that $p_{i}=a_{i}p_{i-1}+p_{i-2}$ and $q_{i}=a_{i}q_{i-1}+q_{i-2}$ with initial conditions $p_{-2}=q_{-1}=0$ and $q_{-2}=p_{-1}=1$.
@@
See [42] and [2, §3.9]. We write the canonical Ostrowski $\alpha$-representation of $n$, starting with the most significant digit, and satisfying the rules (a)–(c) above, as $(n)_{\alpha}$. The inverse mapping, which sends a string $x=e_{t}e_{t-1}\cdots e_{0}$ to the sum $\sum_{0\leq i\leq t}e_{i}q_{i}$, is denoted by $x]_{\alpha}$. The Fibonacci numeration system then is a special case of Ostrowski $\alpha$-numeration, where $\alpha=(3-\sqrt{5})/2$.
In some cases we need to write the $\alpha$-representation of pairs or tuples. We do so by working over a product alphabet, and padding the shorter representation at the front with 0’s, if needed.
@@ 4 Automata, synchronized sequences, and first-order logic
We say a sequence $(s_{n})_{n\geq 0}$ is automatic if there is a deterministic finite automaton with output (DFAO) that takes as input the representation of $n$ in some addable numeration system, and $s_{n}$ is the output associated with the last state reached [2]. Here by ‘addable’, we mean that the addition relation $x+y=z$ is recognizable by an automaton for this numeration system; this is the case for base-$k$ where $k\geq 2$ is an integer, the Fibonacci and Tribonacci numeration systems, and many others. In particular, this is the case for Ostrowski $\alpha$-representation when $\alpha$ is a quadratic irrational [3, 4]. It is clear that such a sequence takes only finitely many values.
We say that a sequence $(s_{n})_{n\geq 0}$ is synchronized if there is a finite automaton that takes the representations of $n$ and $y$ in parallel, and accepts if and only if $y=s_{n}$. In this paper, the particular representations we use are the Ostrowski $\alpha$-representations discussed in the previous section. Moreover, in this paper, we always assume that $n$ and $y$ are represented in the same numeration system.
Synchronized sequences have a number of nice closure properties. These can be found in [46] for $k$-synchronized sequences, but the same proofs also suffice for sequences defined in terms of Ostrowski representations.
@@
    The sequence $(\lfloor(bn+c)/d\rfloor)_{n\geq 1}$ is synchronized for integers $b,c,d$ with $b\geq\max(1,-c)$ and $d\geq 1$.
@@
    If $(f(n))_{n\geq 1}$ and $(g(n))_{n\geq 1}$ are both synchronized then so is their composition $(f(g(n)))_{n\geq 1}$.
@@
There is an algorithm that takes as input a first-order logical formula on natural numbers for $f(n)$, defined in terms of addition, subtraction, multiplication by constants, automatic sequences, synchronized sequences, and computes a synchronized automaton for $f$.
@@
As a consequence of Theorem 2, we get the following corollary. Recall that the subword complexity (aka factor complexity) of a sequence is the function counting the number of distinct contiguous blocks appearing in it.
Corollary 3.
Let ${\bf f}=(f(n))_{n\geq 0}$ be an automatic sequence. Then its subword complexity function $\rho_{\bf f}(n)$ satisfies $\rho_{\bf f}(n)=O(n)$.
Proof.
As explained in [48, §10.5], given an automatic sequence ${\bf f}=(f(n))_{n\geq 0}$ there is a first-order logical formula for the appearance function $A_{\bf f}(n)$, which is the length of the shortest prefix of $\bf f$ containing an occurrence of each of its length-$n$ factors. Hence, by Theorem 2, it follows that there is a synchronized automaton for $A_{\bf f}(n)$. Then from [46, Thm. 8], it follows that $A_{\bf f}(n)=O(n)$. (The proof there technically only was stated for $k$-synchronized sequences, but it goes through in exactly the same way for the more general notion of synchronized sequence over an addable numeration system.) Hence $\rho_{\bf f}(n)$ is also $O(n)$. ∎
@@ 5 The main result and its implementation
We say that a continued fraction is purely periodic of period length $P$ if it is of the form
$a_{1},a_{2},\ldots,a_{P},a_{1},a_{2},\ldots,a_{P},a_{1},a_{2},\ldots,a_{P},\ldots]$ (2)
for some $P\geq 1$.
@@
Suppose $0<\gamma<1$ is a quadratic irrational such that $1/\gamma$ has a purely periodic continued fraction of the form given in Eq. (2). Let $p_{i}/q_{i}$ denote the $i$’th convergent to the continued fraction for $\gamma$. Then for all integers $n\geq 1$ we have
$(n-1)_{\gamma}\,0^{P}]_{\gamma}=q_{P}(n-1)+q_{P-1}\cdot\lfloor n\gamma\rfloor.$
@@
Since $a_{i}=a_{i+P}$ for all $i\geq 1$, we have
$\left[\begin{array}[]{cc}p_{i+P}&p_{i+P-1}\\ q_{i+P}&q_{i+P-1}\end{array}\right]$ $=\left[\begin{array}[]{cc}a_{0}&1\\ 1&0\end{array}\right]\left[\begin{array}[]{cc}a_{1}&1\\ 1&0\end{array}\right]\cdots\left[\begin{array}[]{cc}a_{i+P}&1\\ 1&0\end{array}\right]$
$=\left[\begin{array}[]{cc}a_{0}&1\\ 1&0\end{array}\right]\left[\begin{array}[]{cc}a_{1}&1\\ 1&0\end{array}\right]\cdots\left[\begin{array}[]{cc}a_{P}&1\\ 1&0\end{array}\right]\quad\left[\begin{array}[]{cc}a_{P+1}&1\\ 1&0\end{array}\right]\cdots\left[\begin{array}[]{cc}a_{i+P}&1\\ 1&0\end{array}\right]$
$=\left[\begin{array}[]{cc}a_{0}&1\\ 1&0\end{array}\right]\left[\begin{array}[]{cc}a_{1}&1\\ 1&0\end{array}\right]\cdots\left[\begin{array}[]{cc}a_{P}&1\\ 1&0\end{array}\right]\quad\left[\begin{array}[]{cc}a_{1}&1\\ 1&0\end{array}\right]\cdots\left[\begin{array}[]{cc}a_{i}&1\\ 1&0\end{array}\right]$
$=\left[\begin{array}[]{cc}p_{P}&p_{P-1}\\ q_{P}&q_{P-1}\end{array}\right]\left[\begin{array}[]{cc}q_{i}&q_{i-1}\\ p_{i}&p_{i-1}\end{array}\right],$
@@
$q_{i+P}=q_{P}q_{i}+q_{P-1}p_{i}.$ (4)
Now suppose the Ostrowski $\gamma$-representation of n1n-1 is $\sum_{0\leq i\leq t}e_{i}q_{i}$. Then from Eq. (4), we get
$[(n-1)_{\gamma}0^{P}]_{\gamma}$ $=\sum_{0\leq i\leq t}e_{i}q_{i+P}$
$=\sum_{0\leq i\leq t}e_{i}(q_{P}q_{i}+q_{P-1}p_{i})$
$=\left(q_{P}\sum_{0\leq i\leq t}e_{i}q_{i}\right)+\left(q_{P-1}\sum_{0\leq i\leq t}e_{i}p_{i}\right)$
$=q_{P}(n-1)+q_{P-1}\lfloor n\gamma\rfloor,$
@@
Suppose $0<\gamma<1$ is a quadratic irrational such that $1/\gamma$ has a purely periodic continued fraction of the form given in Eq. (2). Let $p_{i}/q_{i}$ denote the $i$’th convergent to the continued fraction for $\gamma$. Then the sequence $(\lfloor n\gamma\rfloor)_{n\geq 1}$ is $\gamma$-Ostrowski synchronized.
@@
It is easy to see that for each $m\geq 1$ there is a finite automaton ${\tt shift}_{P}$ accepting $0^{P}y$ and $y0^{P}$ in parallel. We sketch the details. The states are $P$-tuples of integers appearing in the Ostrowski representation of $\mathbb{N}$, and the initial state and unique final state is $0^{P}$. Here the “meaning” of a state named $f_{1}\cdots f_{P}$ is that these are the last $P$ symbols of $y$ that the automaton has read. The transition function $\delta$ is defined by $\delta(f_{1}f_{2}\cdots f_{P},[f_{1},g])=f_{2}\cdots f_{P}g$.
Since $q_{P}$ and $q_{P-1}$ are constants, the first-order formula
$\exists u,v\ n=u+1\,\wedge\,{\tt shift}_{P}(u,v)\,\wedge\,v=q_{P-1}\cdot z+q_{P}\cdot u$
@@
Remark 8.
In practice, the automaton for ${\tt shift}_{P}(u,v)$ can sometimes be more easily constructed as follows. We construct a single automaton for ${\tt shift}_{1}(u,v)$, and then create the shifters for $P\geq 2$ by iterating the construction, using a sequence of first-order formulas:
${\tt shift}_{P+1}(u,v):=\exists x\ {\tt shift}_{1}(u,x)\ \wedge\ {\tt shift}_{P}(x,v).$
@@
Suppose $0<\gamma<1$ is a quadratic irrational such that
$\gamma=[0,a_{1},a_{2},\ldots,a_{P},a_{1},a_{2},\ldots,a_{P},a_{1},a_{2},\ldots,a_{P},\ldots],$
and suppose $\alpha,\beta\in\mathbb{Q}(\gamma)$ such that $\alpha\geq 0$ and $\alpha+\beta\geq 0$. Then $(\lfloor n\alpha+\beta\rfloor)_{n\geq 1}$ is $\gamma$-Ostrowski synchronized.
@@
Let $0<\gamma<1$ be a quadratic irrational such that $1/\gamma$ has a purely periodic continued fraction, and let $\alpha,\beta\in\mathbb{Q}(\gamma)$. The first-order theory of the natural numbers with addition and the function $n\rightarrow\lfloor n\alpha+\beta\rfloor$ is decidable.